Hello out there,I'm having some trouble with some problems that I need help with,the first one is,A rectangle has side length 30 mm and diagonal length 34 mm.Find the length of the other side of the rectangle to the nearest whole number.Second one is,Estimate the value of the square root:0.85.Next one is,If x=0.04,what is the value of x?The last one is,Express 0.0049 in scientific notation.Thanks for whoever can help.
Answer by checkley75(3666)

30^2+X^2=34^6 THIS IS A RIGHT TRIANGLE WITH THE FORMULA FOR THE SIDES OF
A^2+B^2=C^2
30^2+X^2=34^2
900+X^2=1190
X^2=1190-900
X^2=290
X=SQRT290
X=17.03 FOR THE OTHER SIDE.
----------------------------------------------------------
SQRT.85~.92
---------------------------------------------------------
X=.04 OR X=4/100 OR X=4% OR X=1/25
----------------------------------------------------------
.0049=4.9*10^-3 BECAUSE WE HAVE MOVED THE DECIMAL 3 PLACES TO THE RIGHT OR WE'VE MULTIPLIED THE NUMBER BY 1,000